It takes me 3 hours to mow my lawn. It takes John 2 hours to mow my lawn. If we mow my lawn together how long will it take us? My first instinct was to add 3 and 2 to get 5 and then divide that by two to find the average and divide that number by two again since they are working on it together. I'm not sure if that is the right thing to do though.
Answer by stanbon(75887)

It takes me 3 hours to mow my lawn. It takes John 2 hours to mow my lawn. If we mow my lawn together how long will it take us?
----
Your DATA:
time = 3 hrs/job ; rate = 1/3 job/hr
--------------------------------------
John DATA:
time = 2 hrs/job ; rate = 1/2 job/hr
--------------------------------------
Together DATA:
time = x hrs/job ; rate = 1/x job/hr
-----------------------------------------
Equation to find "x":
rate + rate = together rate
1/3 + 1/2 = 1/x
(5/6) = 1/x
x = 6/5 hr = 1 hr. and 12 minutes
That is the time if you work together.
=========================================
Cheers,
Stan H.